The Everton College store paid ​$19081908 for an order of 5151 calculators. The store paid ​$88 for each scientific calculator. The​ others, all graphing​ calculators, cost the store ​$5858 each. How many of each type of calculator was​ ordered?

Answer by Boreal(15235)

S=scientific
G=graphic
S+G=51; G=51-S
88S+(51-S)58=1908
88S+2958-58S=1908
This won't solve with a positive number since 30S=-1050
This is because 51 calculators and an order of $1908 is an average price of about $40 per calculator. Neither type is that cheap, so the order for 51 can't work for positive numbers.